OpenStudy (anonymous):

a telephone pole is 10 meters tall and cast a shadow 3 meters long . At the same time, a road sighn, which is 4 meters tall, cast a shadow.How long is the road sighn.

jimthompson5910 (jim_thompson5910):

solve for x 10/3 = 4/x

OpenStudy (anonymous):

10/3 = 4x 3.33x = 4 x = 4/3.33 x=1.2m

OpenStudy (anonymous):

thx